### v2

# Formation of Bound States in Quintessence Alternative Theories

Abstract
We study the formation and behaviour of bound states formed outside the horizon of a black hole in the presence of quintessence matter. Calculating the Regge and Wheeler potential for general metric function, we find that the presence of quintessence influences significantly the metric function and the Hawking temperature. We show that large black holes radiate less in the presence of quintessence matter and it seems to live longer, while small black holes radiate more in comparison with the model in the absence of quintessence. Bound states emerge at large enough quintessence parameter $|w|$ or angular momentum.
I Introduction
Bound states for particles around black holes is a very interesting effect and a detailed study was carried out in Grain:2007gn . A similar study can be done by the investigation of QNMs. In both cases the goal is to find the characteristic complex frequencies for a scalar field which propagates in a black hole background. The constraints which the scalar field has to satisfy are boundary conditions at the black hole horizon and at spatial infinity. However, the physical meanings of bound states and quasi-normal modes are very different and correspond to different energy ranges. In Grain:2007gn a systematic study was carried out of quanta trapped between the finite adS black hole potential barrier and the infinitely thick well, which works as a barrier preventing the particles from reaching infinity. These bound states may have cosmological consequences related to the primordial power spectrum barrau3 ; Carr:2005zd . Also the phenomenology of these bound states and their spectra should be studied in detail and experiments associated with light black holes should also take into account those states.
One basic tool for the examination of the stability of a black hole is to study the Regge and Wheeler potential. Regge and Wheeler regwh first described a decomposition of perturbations on a spherically symmetric background geometry into tensor harmonics and wrote the perturbation equations in terms of these harmonics. They found that it is possible to put the equations for the oscillatory motion into the form of a single second-order Klein-Gordon-type differential equation with an effective potential. Then the study of these equations can give us important information on the stability of the black hole. The first metric that was studied by Regge and Wheeler was the Schwarzschild metric.
The formation and propagation of quantum bound states in the vicinity of a Galileon black hole was discussed in Koutsoumbas:2018gbd . The Galileon black holes are generated by a scalar field coupled to the Curvature Kolyvaris:2011fk ; Rinaldi:2012vy ; Kolyvaris:2013zfa and they are spherically symmetric hairy black holes. Classically this coupling has been studied extensively; it strongly influences the propagation of the scalar field compared to a scalar field minimally coupled to gravity. Also the stability of the background Galileon black hole was studied and it was found that bound states were trapped in a Regge and Wheeler potential well or penetrating the horizon of the Galileon black hole depending on the strength of the derivative coupling of a scalar field to curvature. The formation of quantum bound states requires that there is a radial potential containing a local well for given ranges of black hole parameters in the KG equation.
Recent astronomical observations suggest that our Universe is expanding at an accelerating rate Riess1 ; Perlmutter1999 . Responsible for this expansion is the dark sector of our Universe consisting mainly of dark energy, while the presence of a cosmological constant cannot explain this expansion Carroll2001 ; Copeland2006 . Approximately 70% of the energy density in the universe is in the form of an exotic, negative pressure component, dubbed dark energy. The observational bounds on the properties of the dark energy have continued to tighten. Taking $w$ to be the ratio of pressure to density for the dark energy $w=p_{DE}/\rho_{DE},$ recent observational constraints the range of $1\leq w\leq-\frac{1}{3}$, which causes the acceleration. The border case of w=1w=-1 of the extraordinary quintessence covers the cosmological constant term. As it was recognized Suss ; B ; W , the acceleration represents a challenge for the consistent theory of quantum gravity, because of the outer horizon, which does not allow one to introduce the observable S-matrix in terms of asymptotic past and future states. The outer horizon of de Sitter space significantly differs from the inner horizon of black hole.
The presence of a quintessence field to explain the Universe expansion modifies General Relativity and new modified Gravity theories were proposed with the presence of dynamical scalar fields Hellerman2001 . In these theories black hole solutions were found in the presence of the quintessence matter field. One of the first black hole solutions in the presence of quintessence was discussed in Kiselev:2002dx . New static spherically-symmetric exact solutions of Einstein equations with the quintessential matter surrounding a black hole charged or uncharged were presented. A condition of additivity and linearity in the energy-momentum tensor was found which allows to get correct limits to the known solutions for the electromagnetic static field. Further studies in the presence of quintessence matter were carried out, new black hole solutions were obtained, and the impacts of the quintessence field on their thermal quantities, shadows, and quasinormal modes were discussed Chen2008 -BB5 . The presence of dark matter and quintessence fields outside the horizon of a black hole has an impact on their thermodynamic properties, such as the temperature, and introduces new critical phenomena. These modifications could lead to generalizations of classical thermodynamic results making this area of study very interesting leading to a connection between a black hole event horizon and its entropy Bekenstein1973 ; Hawking1974 .
Motivated by the work in Kiselev:2002dx in which quintessential matter surrounding a background spherically-symmetric black hole backreacts with it dressing it with quintessential hair, we study the formation and behaviour of bound states formed outside the horizon of this hairy black hole. Then considering a general form of the metric function we derive a Schrödinger equation in which the Regge and Wheeler potential appears. Considering the thermodynamics of the solution we discuss the formation and the behaviour of the bound states formed for various values of the quintessence parameter $w$ for the considered hairy black hole. We also consider matter which does not back-react with the quintessence black hole and we study whether the Regge and Wheeler potential is affected by this matter outside the quintessence black hole.
In summary, we study the formation and the behaviour of bound states outside the horizon of a background spherically-symmetric black hole, which is dressed with quintessential hair. Calculating the Regge and Wheeler potential we study the bound states which are trapped outside the horizon of the black hole depending on the quintessence parameter $w.$ We find that the Regge and Wheeler potential for small $|w|$ does not support bound states. However, as the quintessence parameter $|w|$ increases, formation of bound states becomes possible. In addition, when the angular momentum parameter $l$ increases, more bound states are possible and their imaginary parts are much smaller. We also find that, as the charge $Q$ of the black holes increases, the black hole temperature approaches zero on the horizon, in contrast with the case with $Q=0;$ in addition, less bound states are formed. We work with the black hole presented in Kiselev:2002dx in which quintessential matter surrounding a background spherically-symmetric black hole back-reacts dressing it with quintessential hair. We also discuss the thermodynamic properties of the bound states formed for various values of the quintessence parameter $w.$
The work is organized as follows. In section II we review the Regge-Wheeler potential for a fairly general metric function. In section III the quintessence is introduced and some aspects of its thermodynamics are described. Section IV deals with the Regge-Wheeler potential for scalar excitation, while in section V mass eigenstates are calculated. Section VI contains our conclusions.
II Generalized Regge-Wheeler Potential
In this Section we will consider a fairly general metric and we will calculate the Regge-Wheeler potential, following calza . We consider the metric:
$ds^{2}=-G(r)dt^{2}+\frac{dr^{2}}{F(r)}+H(r)d\Omega^{2}~.$ (1)
Then particles with spin $s$ satisfy equation:
- $\frac{H}{G}\partial_{tt}\Psi_{s}(t,r,\theta,\phi)+s\sqrt{\frac{F}{G}}\left(H\frac{G^{\prime}}{G}-H^{\prime}\right)\partial_{t}\Psi_{s}(t,r,\theta,\phi)+FH\partial_{rr}\Psi_{s}(t,r,\theta,\phi)$ (2)
$+\left(\frac{F^{\prime}H}{2}+\left(s+\frac{1}{2}\right)\frac{FG^{\prime}H}{G}+(s+1)FH^{\prime}\right)\partial_{r}\Psi_{s}(t,r,\theta,\phi)$
$+\left[\frac{1}{\sin\theta}\partial_{\theta}(\sin\theta\partial_{\theta})+\frac{2is\cot\theta}{\sin\theta}\partial_{\phi}+\frac{1}{\sin^{2}\theta}\partial_{\phi\phi}-s-s^{2}\cot^{2}\theta\right]\Psi_{s}(t,r,\theta,\phi)$
$+\left[s\frac{FG^{\prime\prime}H}{G}+\frac{3s-2s^{2}}{2}FH^{\prime\prime}-\frac{s}{2}\frac{FG^{\prime 2}H}{G^{2}}+\frac{2s^{2}-s}{4}\frac{FH^{\prime 2}}{H}+\frac{s}{2}\frac{F^{\prime}G^{\prime}H}{G}\right$.
$\left.+\frac{3s-2s^{2}}{2}F^{\prime}H^{\prime}+\frac{2s^{2}+5s}{4}\frac{FG^{\prime}H^{\prime}}{G}\right]\Psi_{s}(t,r,\theta,\phi)=0~.$
We decompose the various fields into radial, angular and time parts, using the ansatz
$\Psi_{s}(t,r,\theta,\phi)=R_{s}(r){\cal Y}_{lm}^{s}(\theta,\phi)e^{-i\omega t}~,$ (3)
where ${\cal Y}_{lm}^{s}(\theta,\phi)$ are the generalized spherical harmonics with spin $s,$ which satisfy the equation
$\left[\frac{1}{\sin\theta}\partial_{\theta}(\sin\theta\partial_{\theta})+\frac{2is\cot\theta}{\sin\theta}\partial_{\phi}+\frac{1}{\sin^{2}\theta}\partial_{\phi\phi}-s-s^{2}\cot^{2}\theta+\lambda_{l}^{s}\right]{\cal Y}_{lm}^{s}(\theta,\phi)=0~,\ \lambda_{l}^{s}=l(l+1)-s(s+1)~.$ (4)
The radial part of the wavefunction will then satisfy:
$A_{s}(B_{s}R_{s}^{\prime})^{\prime}+\left(\frac{\omega^{2}H}{G}+i\omega s\sqrt{\frac{F}{G}}\left(H\frac{G^{\prime}}{G}-H^{\prime}\right)+C_{s}\right)R_{s}=0,\ A_{s}=\sqrt{\frac{F}{G}}\frac{1}{(GH)^{s}},\ B_{s}=H\sqrt{FG}(GH)^{s}~,$ (5)
$C_{s}$ $=$ $s\frac{FG^{\prime\prime}H}{G}+\frac{3s-2s^{2}}{2}FH^{\prime\prime}-\frac{s}{2}\frac{FG^{\prime 2}H}{G^{2}}$ (6)
$+\frac{2s^{2}-s}{4}\frac{FH^{\prime 2}}{H}+\frac{s}{2}\frac{F^{\prime}G^{\prime}H}{G}+\frac{3s-2s^{2}}{2}F^{\prime}H^{\prime}+\frac{2s^{2}+5s}{4}\frac{FG^{\prime}H^{\prime}}{G}-\lambda_{l}^{s}-2s~.$
We will define
$R_{s}\equiv\frac{u_{s}}{\sqrt{H}}~,$
and then the radial part of the equation becomes:
$A_{s}(B_{s}R_{s}^{\prime})^{\prime}$ $=$ $\frac{\sqrt{H}}{G}\left[FGu_{s}^{\prime\prime}+\frac{1}{2}\frac{(FG)^{\prime}u_{s}^{\prime}}{G}\right]$ (7)
$+\frac{1}{4}\left[\frac{FH^{\prime 2}}{H}-F^{\prime}H^{\prime}-\frac{FG^{\prime}H^{\prime}}{G}-2FH^{\prime\prime}\right]\frac{u_{s}}{\sqrt{H}}~.$
Using the tortoise coordinate
$\frac{d}{dr}=\frac{1}{\sqrt{FG}}\frac{d}{dr_{*}}~,$ (8)
we will get
$\frac{du_{s}}{dr}=\frac{1}{\sqrt{FG}}\frac{du_{s}}{dr_{*}},\ \frac{d^{2}u_{s}}{dr^{2}}=\frac{1}{FG}\frac{d^{2}u_{s}}{dr_{*}^{2}}-\frac{(FG)^{\prime}}{2(FG)^{3/2}}\frac{du_{s}}{dr_{*}}~,$ (9)
so that
$A_{s}(B_{s}R_{s}^{\prime})^{\prime}=\frac{\sqrt{H}}{G}\frac{d^{2}u_{s}}{dr_{*}^{2}}+\frac{1}{4}\left[\frac{FH^{\prime 2}}{H}-F^{\prime}H^{\prime}-\frac{FG^{\prime}H^{\prime}}{G}-2FH^{\prime\prime}\right]\frac{u_{s}}{\sqrt{H}}~,$ (10)
and the radial equation takes a Schrödinger-like form.
$-\frac{d^{2}u_{s}}{dr_{*}^{2}}+V_{s}u_{s}=\omega^{2}u_{s}~,$ (11)
where the potentials for $s=0,1,2,\frac{1}{2}$ read
$V_{0}=\nu_{l}^{0}\frac{G}{H}+\frac{1}{\sqrt{H}}\frac{d^{2}}{dr_{*}^{2}}\sqrt{H}=\nu_{l}^{0}\frac{G}{H}+\frac{1}{\sqrt{H}}\left[FG\frac{d^{2}\sqrt{H}}{dr^{2}}+\frac{1}{2}\frac{d(FG)}{dr}\frac{d\sqrt{H}}{dr}\right]~,$ (12)
$V_{1}=\nu_{l}^{1}\frac{G}{H}~,$ (13)
$V_{2}$ $=$ $\nu_{l}^{2}\frac{G}{H}+\frac{1}{2H^{2}}\left(\frac{dH}{dr_{*}}\right)^{2}-\frac{1}{\sqrt{H}}\frac{d^{2}}{dr_{*}^{2}}\sqrt{H}=\nu_{l}^{2}\frac{G}{H}+\frac{FG}{2H^{2}}\left(\frac{dH}{dr}\right)^{2}$ (14)
$-\frac{1}{\sqrt{H}}\left[FG\frac{d^{2}\sqrt{H}}{dr^{2}}+\frac{1}{2}\frac{d(FG)}{dr}\frac{d\sqrt{H}}{dr}\right]~,$
$V_{\frac{1}{2}}=\nu_{l}^{1/2}\frac{G}{H}\pm\sqrt{\nu_{l}^{1/2}}\frac{d}{dr_{*}}\sqrt{\frac{G}{H}}=\nu_{l}^{1/2}\frac{G}{H}\pm\sqrt{\nu_{l}^{1/2}FG}\frac{d}{dr}\sqrt{\frac{G}{H}}~,\nu_{l}^{s}=l(l+1)-s(s-1).$ (15)
In the case, where (F=G=f, $H=r^{2}$) with the metric
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2}~,$ (16)
the potentials become
$V_{0}=f\left[\frac{l(l+1)}{r^{2}}+\frac{f^{\prime}}{r}\right]~,\ V_{1}=l(l+1)\frac{f}{r^{2}}~,\ V_{2}=f\left[\frac{l(l+1)-2}{r^{2}}-\frac{2f}{r^{2}}-\frac{1}{r}\frac{df}{dr}\right]~,$ (17)
$V_{\frac{1}{2}}=f\left[\frac{l(l+1)+\frac{1}{4}}{r^{2}}\pm\sqrt{l(l+1)+\frac{1}{4}}\frac{d}{dr}\sqrt{\frac{f}{r^{2}}}\right]~.$ (18)
As we can see the spin changes the structure of the Regge-Wheeler potential. In this work we concentrate on the $s=0$ case.
III Quintessence Matter and its Thermodynamics
In this section we will introduce quintessence matter outside the horizon of a black hole, we study the thermodynamics of the coupled system and discuss the formation of bound states.
III.1 Description of the Quintessence
As we have already discussed, quintessence is a scalar field, whose equation of state parameter $w$ is defined as the ratio of its pressure $p$ and its energy density $\rho$ and is given by
$w=\frac{p}{\rho},\ \ -1\leq w\leq-\frac{1}{3}~.$ (19)
Writing the line element in the form:
$ds^{2}=(1+h)dt^{2}-\frac{dr^{2}}{1+h}-r^{2}d\Omega^{2}=fdt^{2}-\frac{dr^{2}}{f}-r^{2}d\Omega^{2},\ f\equiv 1+h,$ (20)
the components of the energy-momentum tensor Kiselev:2002dx read:
$T_{t}^{t}=-\frac{1}{2r^{2}}(h+rh^{\prime})=\rho,\ T_{\theta}^{\theta}=T_{\phi}^{\phi}=-\frac{1}{4r}(2h^{\prime}+rh^{\prime\prime})=-\frac{1}{2}\rho(3w+1)~.$ (21)
There exist two solutions for the Einstein equations, the well-known Schwarzschild solution $h=-\frac{2M}{r}$ and a novel one
$h=\frac{a}{r^{1+3w}},$ (22)
where $a$ and $w$ are constants. The Einstein equations yield:
$\rho=\frac{a}{2}\frac{3w}{r^{3(w+1)}},\ R=3aw\frac{1-3w}{r^{3(1+w)}}~.$ (23)
Notice that $R$ is singular at $r\to 0,$ except when w=0,13,1.w=0,\ $\frac${1}{3},\ -1.
Additivity and linearity yield the final solution for the metric: $f=1-\frac{2M}{r}-\frac{a}{r^{1+3w}}.$ The generalized solution to the Reissner–Nordström–anti-deSitter black hole reads:
$f=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+\frac{r^{2}}{L^{2}}-\frac{a}{r^{1+3w}}~.$ (24)
The metric function in the AdS Black Hole, apart from the usual dependence mass and charge, it also depends on two more parameters $a$ and $w,$ which play a crucial role in the thermodynamics of the system and the formation of bound states.
III.2 Thermodynamics in AdS Spacetime
The mass can be expressed in terms of the remaining parameters and the horizon, since $f(r_{+})=0:$
$M=\frac{r_{+}}{2}\left[1+\frac{Q^{2}}{r_{+}^{2}}+\frac{r_{+}^{2}}{L^{2}}-\frac{a}{r_{+}^{1+3w}}\right]~.$ (25)
The Hawking temperature is given by:
$T=\frac{1}{4\pi}f^{\prime}(r_{+})=\frac{1}{4\pi}\left[\frac{2M}{r_{+}^{2}}-\frac{2Q^{2}}{r_{+}^{3}}+\frac{2r_{+}}{L^{2}}+\frac{a(1+3w)}{r_{+}^{2+3w}}\right]~,$ (26)
and substituting the mass M it becomes:
$T=\frac{\frac{3r_{+}}{L^{2}}-\frac{Q^{2}}{r_{+}^{3}}+\frac{1}{r_{+}^{2}}(r_{+}+3awr_{+}^{-3w})}{4\pi}~,$ (27)
while the entropy is given by
$S=\int\frac{dM}{T}=\frac{A}{4}=\pi r_{+}^{2}~.$ (28)
In the following we will investigate how the above quantities are affected by the quintessence parameters $a$ and $w.$ We set $L=1.$
In Fig. 1, upper left panel, we plot the metric function for Q=0.00,w=0.80Q=0.00,\ w=-0.80 and various values of the $a$ parameter. The metric tends to take smaller values as $a$ increases. In the upper right panel of the same figure, we plot the metric function for Q=0.00,a=0.80Q=0.00,\ a=0.80 and various values of the $w$ parameter. There is a different effect of $w$ on the small $(r\leq 1.0)$ and the large black holes. In the panel down left, we plot the metric function for Q=0.20,w=0.80Q=0.20,\ w=-0.80 and various values of the $a$ parameter. Finally, in the panel down right, we plot the metric function for Q=0.20,a=0.80Q=0.20,\ a=0.80 and various values of the $w$ parameter. As we can see, the existence of a non-zero charge $Q$ in the black hole tends to decrease the value of the metric function and to increase the $r_{+}$ value. The qualitative characteristics of the graphs are similar to the $Q=0.00$ cases.
Figure 1: (a) Metric function $g\equiv f(r)$ versus $r$ for $Q=0.00,$ w=0.80w=-0.80 and $a=0.20,$ (blue curve, highest) $a=0.40$ (purple curve) $a=0.60$ (brown curve) and $a=0.80$ (cyan curve, lowest). (b) Metric function $g\equiv f(r)$ versus $r$ for $Q=0.00,$ $a=0.80$ and w=0.40,w=-0.40, (blue curve, highest on the right part) w=0.60w=-0.60 (purple curve) w=0.80w=-0.80 (brown curve) and w=1.00w=-1.00 (cyan curve, lowest on the right part). (c) Metric function $g\equiv f(r)$ versus $r$ for $Q=0.20,$ w=0.80w=-0.80 and $a=0.20,$ (blue curve, highest) $a=0.40$ (purple curve) $a=0.60$ (brown curve) and $a=0.80$ (cyan curve,lowest). (d) Metric function $g\equiv f(r)$ versus $r$ for $Q=0.20,$ $a=0.80$ and w=0.40,w=-0.40, (blue curve, highest on the right part) w=0.60w=-0.60 (purple curve) w=0.80w=-0.80 (brown curve) and w=1.00w=-1.00 (cyan curve, lowest on the right part).
In Fig. 2, upper left panel, we plot the temperature versus $r_{h}$ for Q=0.00,w=0.80Q=0.00,\ w=-0.80 and various values of the $a$ parameter. The temperature decreases for increasing $a.$ Moreover the temperature is not a monotonic function and the small and large black holes respond differently to the quintessence. In the upper right panel of the same figure, we plot the temperature versus $r_{h}$ for Q=0.00,a=0.80Q=0.00,\ a=0.80 and various values of the $w$ parameter. As $|w|$ increases, the temperature is smaller for large $r_{h},$ while for modest $r_{h}$ it is larger. In the panel down left, we plot the temperature versus $r_{h}$ for Q=0.20,w=0.80Q=0.20,\ w=-0.80 and various values of the $a$ parameter. Finally, in the panel down right, we plot the temperature versus $r_{+}$ for Q=0.20,a=0.80Q=0.20,\ a=0.80 and various values of the $w$ parameter. For $Q=0.00,$ at some horizon size, the temperature has a minimum. The temperature of the black hole becomes smaller as $|w|$ increases from $\frac{1}{3}$ to $1$ for large enough black holes (big $r_{+}$), while it becomes larger for small $r_{+}.$ Thus large black holes radiate less for large $|w|,$ while small black holes radiate more. Thus large black holes seem to be longer-lived for large $|w|.$ For $Q>0$ the new feature is that the temperature presents a local minimum for small values of $r_{h},$ while it had a maximum for $Q=0.$
Figure 2: (a) Temperature versus $r_{h}\equiv r_{+}$ for $Q=0.00,$ w=0.80w=-0.80 and $a=0.20,$ (blue curve, highest) $a=0.40$ (purple curve) $a=0.60$ (brown curve) and $a=0.80$ (cyan curve, lowest). (b) Temperature versus $r_{h}\equiv r_{+}$ for $Q=0.00,$ $a=0.80$ and w=0.40,w=-0.40, (blue curve, highest on the right part) w=0.60w=-0.60 (purple curve) w=0.80w=-0.80 (brown curve) and w=1.00w=-1.00 (cyan curve, lowest on the right part). (c) Temperature versus $r_{h}\equiv r_{+}$ for $Q=0.20,$ w=0.80w=-0.80 and $a=0.20,$ (blue curve, highest) $a=0.40$ (purple curve) $a=0.60$ (brown curve) and $a=0.80$ (cyan curve, lowest). (d) Temperature versus $r_{h}\equiv r_{+}$ for $Q=0.20,$ $a=0.80$ and w=0.40,w=-0.40, (blue curve, highest on the right part) w=0.60w=-0.60 (purple curve) w=0.80w=-0.80 (brown curve) and w=1.00w=-1.00 (cyan curve, lowest on the right part).
IV Regge-Wheeler potentials
We first consider the metric in the AdS spacetime with quintessence
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2},\ \ \ f(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+r^{2}-\frac{a}{r^{1+3w}}~.$ (29)
We concentrate specifically on the potential for scalar perturbations. The potential $V_{0}$ for zero spin (scalar) perturbations reads
$V_{0}=f(r)g_{0}(r),\ g_{0}(r)=\frac{f^{\prime}(r)}{r}+\frac{l(l+1)}{r^{2}}=2-\frac{2Q^{2}}{r^{4}}+\frac{2M}{r^{3}}+\frac{l(l+1)}{r^{2}}+\frac{a(1+3w)}{r^{3+3w}}~.$ (30)
We start in Fig. 3, left panel, with the Regge-Wheeler potentials, when $a=0,$ so that quintessence plays no role at all; we set $M=0.05$ and let $Q$ take the values 0.00, 0.04, 0.070.00,\ 0.04,\ 0.07 and $0.10.$ It is obvious that, as $Q$ increases, the left barrier of the potential becomes smaller, and at $Q=0.10$ it disappears completely. In this case it is obvious that the potential cannot support bound states: any scalar excitation with $l=0$ will end into the black hole horizon. This characteristic is independent from the role of quintessence. We continue in the right panel with the case a=0.8,w=0.80a=0.8,\ w=-0.80 and $M=0.05$ and the same values for the charge $Q$ as before. The picture is qualitatively the same, but the presence of the quintessence tends to reduce the size of the potentials.
Figure 3: Left panel: Potentials for a=0,l=0,M=0.05a=0,\ l=0,\ M=0.05 and four values of the charge: $Q=0.00$ (pink curve, highest), $Q=0.04$ (blue curve), $Q=0.07$ (khaki curve) and $Q=0.10$ (green curve, lowest). Right panel: Potentials for a=0.8,w=0.80,M=0.05a=0.8,\ w=-0.80,\ M=0.05 and $Q=0.00$ (pink curve, highest), $Q=0.04$ (blue curve), $Q=0.07$ (khaki curve) and $Q=0.10$ (green curve, lowest).
We go ahead in Fig. 4 with the potentials at four different values of $w,$ namely 0.40,0.60,0.80,1.00,-0.40,\ -0.60,\ -0.80,\ -1.00, while l=0,M=0.05,l=0,\ M=0.05, for $Q=0.00$ (left panel) and $Q=0.04$ (right panel). We observe that, for both values of the charge, the potential for w=0.40w=-0.40 does not support bound states, but a left barrier gradually develops, as $|w|$ increases and bound states become possible.
Figure 4: Left panel: Potentials for a=0.8,l=0,M=0.05,Q=0.00,a=0.8,\ l=0,\ M=0.05,\ Q=0.00, and four values for the parameter $w:$ w=0.40w=-0.40 (blue curve, no left barrier), w=0.60w=-0.60 (pink curve), w=0.80w=-0.80 (khaki curve) and w=1.00w=-1.00 (green curve, highest left barrier). Right panel: Same as in the left panel, but $Q=0.04.$
Fig. 5, left panel, depicts potentials for M=0.05,Q=0.00,w=0.60M=0.05,\ Q=0.00,\ w=-0.60 and l=0, 1, 2.l=0,\ 1,\ 2.. It clearly shows that the presence of angular momentum builds gradually a barrier on the left, so that more bound states may be formed. On the other hand, in the right panel one may inspect the potentials for M=0.05,Q=0.00,l=0,w=0.60M=0.05,\ Q=0.00,\ l=0,\ w=-0.60 and a=0.2, 0.4, 0.6, 0.8.a=0.2,\ 0.4,\ 0.6,\ 0.8. The potential decreases for increasing $a.$
Figure 5: Left panel: Potentials for a=0.8,M=0.05,Q=0.00,w=0.60a=0.8,\ M=0.05,\ Q=0.00,\ w=-0.60 and three values for the angular momentum: $l=0$ (blue curve, lowest), $l=1$ (pink curve), and $l=2$ (green curve, highest). Right panel: Potentials for l=0,M=0.05,Q=0.00,w=0.60l=0,\ M=0.05,\ Q=0.00,\ w=-0.60 and four values for the parameter $a:$ $a=0.2$ (blue curve, highest), $a=0.4$ (pink curve), $a=0.6$ (khaki curve), and $a=0.8$ (green curve, lowest).
V Calculation of Bound States
We now proceed with the calculation of the masses of the bound states, along with their imaginary parts. The latter signal the instability of these states, since the latter may tunnel through the barrier on the left and end up inside the horizon. To find the real part of the masses we follow Grain:2007gn and consider the equation:
$\int_{r_{2}}^{r_{3}}\frac{\sqrt{\omega^{2}-V_{0}}}{f(r)}dr=\left(n+\frac{1}{2}\right)\pi,n=0,\ 1,\ 2,\ \dots.$ (31)
and we tune $\omega,$ so that the equation is satisfied for some integer $n.$ The states with $n=0$ are the fundamental massive modes, while the ones with n=1, 2,n=1,\ 2,\ $\dots$ are the overtones. In general a small number of overtones exists, since the energy levels eventually get higher than the left barrier. We also consider various values of the angular momentum quantum number $l;$ we will show the results for $l=0$ and $l=1.$
To determine the imaginary parts of these states, interpreted as bandwidths, we fix the real part of $\omega,$ that we found in the previous step (call it $m$), and we calculate the bandwidth via:
$\gamma=m\frac{\exp\left(-2\int_{r_{1}}^{r_{2}}\frac{\sqrt{V_{0}-m^{2}}}{f(r)}dr\right)}{1+\exp\left(-\frac{2\pi p^{2}(r_{max})}{\sqrt{2\frac{d^{2}p^{2}(r_{max})}{dr_{*}^{2}}}}\right)}.$ (32)
Notice that $\frac{d}{dr_{*}}=f(r)\frac{d}{dr},$ so that
$\frac{d^{2}(p^{2})}{dr_{*}^{2}}=f^{2}(r)\frac{d^{2}(p^{2})}{dr^{2}}+f(r)f^{\prime}(r)\frac{d(p^{2})}{dr},\ p^{2}=m^{2}-V_{0}.$ (33)
Notice that $r_{1}$ and $r_{2}$ are the two points common to the horizontal line at $m^{2}$ and the potential $V_{0},$ which lie nearest to the origin, where the $m$ horizontal line cuts the potential curve, while $r_{3}$ is the most distant cut. In addition, $r_{max}$ is the point, where the potential attains its maximal value. After we are done with these calculations, we have the mass eigenvalues in the form $m-i\gamma.$
We now report the mass eigenvalues and their bandwidths for various values of $w$ and $a.$ In Fig. 6, left panel, we depict the dependence of the masses on $w$ for Q=0.00,M=0.05,a=0.8Q=0.00,\ M=0.05,\ a=0.8 and $l=0.$ It appears that there exist no bound states for $|w|$ smaller than some value (about 0.55), while the bandwidths (denoted by error bars) increase as we approach this value. The masses do not change much, unless near the value w=1.w=-1. In the right panel we show the corresponding results for Q=0.00,M=0.05,w=0.6Q=0.00,\ M=0.05,\ w=-0.6 and $l=0$ versus $a.$ There is an almost linear dependence of the masses on $a,$ while the bandwidths are more or less constant. This behaviour must be correlated with the changes in the potentials, brought about by the changes in $w.$ In Fig. 4, left panel, it is seen that, as $|w|$ decreases, the potential gradually loses its left barrier. This is why we are not able to find masses for $|w|$ smaller than some value. In addition, in Fig. 5, right panel, it is seen that the potentials get larger values for small values of $a,$ so one would expect larger values for the masses in this region, which is actually observed in Fig. 6.
Figure 6: Left panel: Masses versus $w$ for Q=0.00,M=0.05,a=0.8Q=0.00,\ M=0.05,\ a=0.8 and $l=0.$ Right panel: Masses versus $a$ for Q=0.00,M=0.05,w=0.6Q=0.00,\ M=0.05,\ w=-0.6 and $l=0.$
In Fig. 7, left panel, we depict the dependence of the masses on $w$ for Q=0.00,M=0.05,a=0.8Q=0.00,\ M=0.05,\ a=0.8 and $l=1.$ This larger angular momentum allows for eigenvalues at values of $|w|$ smaller than before. This is due to the fact that the decrease in the height of the barrier on the left, due to the decrease of $|w|,$ is partially compensated by the increase, due to the larger angular momentum. In addition the bandwidths are very much smaller and this is due to the large barrier on the left due to the angular momentum (compare Fig. 5); this barrier does not allow the particles go into the horizon easily, thus their imaginary parts are reduced to the size of the symbols. The masses are larger than before and the qualitative behaviour versus $w$ is the same. In the right panel we depict the corresponding results for Q=0.00,M=0.05,w=0.6Q=0.00,\ M=0.05,\ w=-0.6 and $l=1$ versus $a.$ The dependence of the masses on $a$ is, once more, linear.
Figure 7: Left panel: Masses versus $w$ for Q=0.00,M=0.05,a=0.8Q=0.00,\ M=0.05,\ a=0.8 and $l=1.$ Right panel: Masses versus $a$ for Q=0.00,M=0.05,w=0.6Q=0.00,\ M=0.05,\ w=-0.6 and $l=1.$
VI Conclusions
We have studied black hole solutions in the presence of quintessence. In particular we concentrated in the bound states of scalar excitations outside the horizon. It has been shown that the parameters of the quintessence influence significantly the metric function and the Hawking temperature. It has turned out (Fig. 2) that, for w=0.40,w=-0.40, large black holes $(r\gtrsim 0.6)$ radiate more than small ones, while, as $|w|$ increases, the behaviour changes: for w=1.00,w=-1.00, small black holes radiate more than big ones. Thus large black holes with large $|w|$ seem to be longer lived, compared to large black holes with small $|w|.$ The situation is reversed for small black holes. In the long run, one may expect to find mainly large black holes with large $|w|,$ or small black holes with small $|w|,$ if other factors (e.g. angular momentum or charge) do not change. In addition, black holes with large $a$ have smaller temperatures and radiate less than the ones with small (or not at all) $a.$
The Regge-Wheeler potentials depend on the charge $Q$ of the black hole: when $Q$ gets large enough, there is no left barrier, so the scalar particles will meet no obstacle to dive into the black hole (Figs. 3 and 4). This has nothing to do with the quintessence. The presence of quintessence enhances the left barrier to smaller values. For $Q$ values modest enough, the presence of the quintessence may induce qualitative changes: the potential for w=0.40w=-0.40 does not support scalar bound states, but larger $|w|$ may create a potential with a barrier on the left, which can support bound states. The barrier on the left may be enhanced by increasing the angular momentum quantum number $l$ (Fig. 5, left panel). This suggests that the scalar particles with large angular momentum are longer lived than the ones with smaller $l.$
The study of the potentials suggests that $|w|$ should be bigger than some bound, if bound states are to be supported, as depicted in Fig. 6, left panel. The presence of quintessence reduces the values of the potentials (Fig. 5, right panel), so we expect that the masses of the possible bound states will decrease with increasing $a.$ The decrease for increasing parameter $a$ is almost linear (right panel of the same figure). Also, the states with large angular momenta $l$ have larger masses and smaller bandwidths than the ones with small $l.$ The relevant results are to be found in Fig. 7. In short, bound states emerge at $a=0,$ while if $a\neq 0,$ bound states are only present for large enough $|w|$ and/or $l.$
As it is well known Chiba:1999wt , the observational evidence for a cosmological constant, which is responsible for the expansion of the Universe, gives the bounds 1<w<1/3.-1<w<-1/3. In accord with this, we have found that bound states may be formed for large absolute values of the quintessence parameter $|w|,$ contributing in this way to the expansion of the Universe.
In our study we have not considered extra matter surrounding the quintessence black hole. If this matter distribution does not back-react on the background metric, the no-hair theorems do not allow to influence the structure of the background black hole. However, if we perturb this matter distribution calculating the QNMs, an observer outside the black hole will detect the properties of the quintessence matter. We leave this study for future work.